OpenStudy (faman39):

Danika has saved enough money to pay cash for a used car. She wants to buy a car that costs $ 4,900 before taxes. If she will be charged 6% sales tax, what is the total amount she will pay for the car?

OpenStudy (anonymous):

Total amount=(6% of 4,900) +4900 =294+4900 =5194

Directrix (directrix):

Total cost = 4900 + 6% of 4900 = 1.06(4900) = $5194 --> same as above

OpenStudy (faman39):

thanks u so much!